# Data-Driven Robust Predictive Control with Interval Matrix Uncertainty Propagation

### 16 Mar 2026 v1

@@ Abstract
This paper presents a new data-driven robust predictive control law, for linear systems affected by unknown-but-bounded process disturbances. A sequence of input-state data is used to construct a suitable uncertainty representation based on interval matrices. Then, the effect of uncertainty along the prediction horizon is bounded through an operator leveraging matrix zonotopes. This yields a tube that is exploited within a variable-horizon optimal control problem, to guarantee robust satisfaction of state and input constraints. The resulting data-driven predictive control scheme is shown to be recursively feasible and practically stable. A numerical example shows that the proposed approach compares favorably to existing methods based on zonotopic tubes and is competitive with an approach combining set-membership system identification and model-based predictive control.
@@ I Introduction
In this paper, a new data-driven robust predictive control scheme is proposed, for linear discrete-time systems with unknown-but-bounded process disturbances. The set of systems compatible with a collection of state-input data is characterized in terms of interval matrices. Then, the uncertainty propagation along the prediction horizon is approximated by using a suitable operator based on matrix zonotopes, similarly to what is done in [3, 25]. The considered data-driven interval matrix uncertainty representation has a twofold benefit: i) it allows one to guarantee recursive feasibility and practical stability of a variable-horizon robust DDPC scheme, without requiring any assumption on the data density; ii) it provides an efficient way to compute a tube bounding the uncertainty propagation, which turns out to be less conservative than that used in [13].
The rest of the paper is organized as follows. Section II introduces some preliminary material on matrix operations and set approximations. The DDPC problem formulation is presented in Section III. The proposed solution based on interval matrix uncertainty propagation is detailed in Section IV, which also states the closed-loop properties of the control scheme. Section V reports numerical results showing the performance of the proposed approach. Finally, some conclusions are drawn in Section VI.
@@ II Notation and preliminaries
For sets of real matrices $\mathcal{A}$, $\mathcal{B}$, the following operations are considered: $\mathcal{A}\oplus\mathcal{B}=\{A+B|\,\,A\in\mathcal{A},B\in\mathcal{B}\}$, $\mathcal{A}\mathcal{B}=\{AB|\,\,A\in\mathcal{A},B\in\mathcal{B}\}$, $\mathcal{A}^{j}=\{A^{j}|\,\,A\in\mathcal{A}\}$. An interval matrix $\mathcal{I}=\{M|\,\,C-\Delta\leq M\leq C+\Delta\}$, where the operator $\leq$ represents elementwise inequality, is compactly denoted as $\mathcal{I}=C\oplus\left\llbracket\Delta\right\rrbracket$. A matrix zonotope $\mathcal{M}=\Big\{M|M=M_{C}+\sum_{i=1}^{n_{g}}G_{i}\beta_{i},\;|\beta_{i}|\leq 1\Big\}$, where $M_{C}$ is the center and $G_{i}$, $i=1,\dots,n_{g}$, are the generators, is denoted as $\mathcal{M}=\langle M_{C};\,G_{1},\ldots,\,G_{n_{g}}\rangle$.
@@
Given a matrix $M\in\mathbb{R}^{l\times q}$, we define the set of matrices $\mathcal{E}(M)=\left\{M^{(k)}\right\}_{k=1}^{lq}$, containing all matrices $M^{(k)}\in\mathbb{R}^{l\times q}$ defined as
$M^{(k)}_{ij}=\begin{cases}M_{ij},&\text{if }k=(i-1)l+j,\\ 0,&\text{otherwise},\end{cases}$
@@
If $j$ successive iterations of the operator $\mathbb{T}_{\mathcal{I}}$ are performed on the zonotope $\mathcal{M}$, one obtains the set approximation
@@ III Problem Formulation
Consider the uncertain linear time-invariant system
@@
where $x(k)\in\mathbb{R}^{n}$ is the system state at time $k$, $u(k)\in\mathbb{R}^{m}$ is the control input, $w(k)\in\mathbb{R}^{n}$ is the process disturbance and A,BA,\,B are unknown system matrices. It is assumed that $w(k)$ is bounded in scaled infinity norm, i.e.
@@
Suppose that input-state data of length $T$ generated by system (5) are available. These samples are organized into the data matrices
$X_{+}=\left[x(1)\;x(2)\;\cdots\;x(T)\right],$
$X_{-}=\left[x(0)\;x(1)\;\cdots\;x(T-1)\right],$
$U_{-}=\left[u(0)\;u(1)\;\cdots\;u(T-1)\right].$
@@
for any possible realization of $w(k)$, with $\mathcal{X}$, $\mathcal{U}$ being convex polytopic sets. To proceed, we first compute data-driven interval matrix bounds for the unknown system matrices $A$ and $B$, that are consistent with the collected data.
@@ III-A Data-driven Interval System Matrices
where $\Delta_{S}=\overline{w}\,\mathbbm{1}^{\top}|\Phi^{\dagger}|$.
@@ III-B Set-membership identification of Interval System Matrices
where $X_{+,i}$ is the $i$-th row of $X_{+}$. Therefore, the feasible set for each $\theta_{i}$ is a bounded polyhedron, and the tightest interval bounds on the entries of $\Theta$ can be computed element-wise by solving $n(n+m)$ linear programs subject to the constraints (18). Once such bounds are derived, an uncertainty model in the form (14) is available. This will be used as a benchmark to evaluate the performance of the proposed DDPC approach.
@@ III-C Robust Predictive Control
IV Data-driven Robust Predictive Control
@@
$\mathcal{B}^{\mathcal{W}}(j)=\sum_{i=0}^{j-1}\mathcal{I}^{\mathcal{W}}(i).$ (30)
@@
[…]
Proof. From (21) one has that $\mathcal{A}_{K}\subseteq\mathcal{I}_{A_{K}}$. Hence, as a consequence of (4) one gets $\mathcal{A}_{K}^{j}\mathcal{I}_{\Delta}\subseteq\mathcal{I}^{\Delta}(j)$ in (29) and $\mathcal{A}_{K}^{j}\mathcal{W}\subseteq\mathcal{I}^{\mathcal{W}}(j)$ in (31). Therefore, the result follows from definitions (27) and (32).∎
@@
Denoting by $N^{*}_{k}$, $\mathbf{v^{*}(k)}$, $\mathbf{z^{*}(k)}$, the optimal solution of problem (34), the nominal control input in (20) is then chosen as
@@
as in standard MPC schemes.
The only difference between problem (34) and problem (24) is that the constraints (34c)-(34e) are more conservative than the corresponding constraints in problem (24), due to the inclusion (33). This implies that the closed-loop system resulting from the application of the DDPC scheme (34)-(35) to system (5) will robustly satisfy the constraints (6), whatever are the system matrices $A,B$ and the process disturbance realization $w(k)\in\mathcal{W}$.
It is worth stressing that, while the sets $\mathcal{B}_{k}^{\Delta}(j)$ depend on the nominal state and input sequences $\mathbf{z}_{k}$ and $\mathbf{v}_{k}$, which are optimization variables of problem (34), the interval matrices $\mathcal{I}^{\Delta}(j)$, $\mathcal{I}^{\mathcal{W}}(j)$ and $\mathcal{B}^{\mathcal{W}}(j)$ can be computed offline for all $j$, since they are independent of $\mathbf{v}_{k}$ and $\mathbf{z}_{k}$. This is a key feature of the proposed approach because it allows to significantly limit the computational burden required for the online solution of problem (34).
In order to derive the closed-loop properties of the DDPC scheme (34)-(35), the following assumptions are enforced.
@@
The terminal set $\mathcal{X}_{f}$ in (34e) is such that $\mathcal{X}_{f}\subseteq\mathcal{X}$ and $K\mathcal{X}_{f}\subseteq\mathcal{U}$. Moreover, it is a robust positive invariant (RPI) set for the uncertain system $x(k+1)=A_{K}x(k)+w(k)$, which means that $A_{K}x+w\in\mathcal{X}_{f}$, $\forall x\in\mathcal{X}_{f}$, $\forall A_{K}\in\mathcal{A}_{K}$ in (21), and $\forall w\in\mathcal{W}$.
@@
Let Assumptions 1-3 be satisfied. Then, the closed-loop trajectories $x(k)$ of system (5), under the control law (20), (34)-(35), reach the set $\mathcal{X}_{f}$ in at most $\lfloor J_{0}^{*}/\gamma\rfloor$+1 steps and converge asymptotically to the set $\mathcal{X}_{\infty}=\sum_{i=0}^{\infty}\mathbb{T}_{\mathcal{I}_{A_{K}}}^{i}(\mathcal{W})$.
@@
It is worth remarking that the advantage of the proposed approach stems from the combined use of two distinct features. First, the zonotopic uncertainty propagation along the prediction horizon, through the operator $\mathbb{T}^{j}$, permits to explicitly incorporate the effect of the data-driven uncertainty model (12)-(13) in the robust optimal control problem (34). Secondly, the adoption of the interval-matrix-based tube $\mathcal{B}_{k}(j)$ has a twofold effect: (i) it is key to prove recursive feasibility of the DDPC scheme; (ii) it allows one to cast the constraints of problem (34) in the computationally lightweight form (37).
The adoption of a variable-horizon predictive control scheme is also instrumental to achieve recursive feasibility and, at the same time, it provides further flexibility to the objective function $J_{k}$ in (25), via the tuning parameter $\gamma$. Summing up, problem (34) is a mixed-integer program with linear constraints. Such a problem can be solved exactly by using standard solvers (see, e.g., [16]), or by adopting efficient heuristic procedures [20, 23].
@@ V Numerical results
In this section, numerical results for a double-integrator example are presented to evaluate the performance of the proposed Data-Driven Interval Matrix Predictive Control (DD-IMPC) approach. This control strategy is compared to the Tube-Based Zonotopic Predictive Control (TZPC) method introduced in [13]. As a further comparison benchmark, we consider the Set-Membership Interval Matrix Predictive Control (SM-IMPC) scheme, that is a version of DD-IMPC, in which the uncertainty model (14) is estimated via the procedure in Section III-B. The simulations are carried out using MATLAB on a laptop equipped with the Apple M4 processor and 16 GB of RAM. The code is written in MATLAB and is publicly available at https://github.com/RenatoQuartullo/Data-Driven-Interval-Matrices-PC.
Consider a discrete-time system (5), in which the true system matrices are
@@
while the input command must satisfy $|u(k)|\leq 2$. The matrix gain $K$ and the terminal set $\mathcal{X}_{f}$ are calculated following the procedures described in [19, Sec. 5.2] and [6, Ch. 10], respectively. In DD-IMPC (and SM-IMPC), the function $\ell(\cdot)$ in the cost $J_{k}$ is quadratic, i.e., $J_{k}=\gamma N_{k}+\sum_{j=0}^{N_{k}-1}\|v_{k}(j)-Kz_{k}(j)\|_{2}^{2}$ with $\gamma=1$. The variable-horizon problem (34) is solved by enumeration for each horizon length $N_{k}\in\{1,\ldots,25\}$. The true system is used only to generate the data matrices $X_{-},\,X_{+},\,U_{-}$, collecting a trajectory of length $T=50$.
In this experiment, we aim to evaluate the feasible domain (FD) of the considered control schemes and their sensitivity to the available data. The FDs are computed as the convex hull of the initial conditions for which the initial optimization problem is feasible. A total of 300 initial conditions are generated by uniformly gridding the admissible state set $\mathcal{X}$. To ensure a fair comparison, the feasible domain of TZPC (which relies on a fixed prediction horizon) is defined as the set of all initial states for which the corresponding optimization problem is feasible for at least one horizon length $N\in\{1,\ldots,25\}$. The FD evaluation is repeated 50 times using independently generated input-state datasets. Table I reports the average area of the FDs, their standard deviation, and the number of runs in which the FD is empty for each method. The results show that DD-IMPC achieves a larger FD on average and exhibits significantly lower sensitivity to the available data compared to TZPC. It is also worth noting that the FD of DD-IMPC is very close to that yielded by SM-IMPC, that, as expected, is essentially unaffected by the dataset. Figure 1 illustrates the FDs obtained in one representative run in which TZPC produces a non-empty feasible domain.
@@
To assess the sensitivity of the considered control schemes to the size of the process disturbance, we consider $\mathcal{W}=\epsilon_{W}\left\llbracket\overline{w}\right\rrbracket$, with $\epsilon_{W}\in\{0.2,0.4,\ldots,0.8,2\}$. For each value of $\epsilon_{W}$, the experiment is repeated 50 times using independently generated datasets. The average area of the resulting FDs is reported in Fig. 2 (solid lines), along with the minimum and maximum values across the runs (dashed lines). It can be observed that, on average, DD-IMPC yields a larger feasible domain than TZPC. Moreover, TZPC exhibits a strong sensitivity to $\epsilon_{W}$, as the area of the FD drops significantly for larger values of the disturbance bound. Moreover, the low data sensitivity of DD-IMPC is confirmed by the small gap between the minimum and maximum area profiles, whereas this effect is much more pronounced for TZPC. In addition, for $\epsilon_{W}\geq 0.8$, TZPC results in an empty feasible region in some runs, while for $\epsilon_{W}\geq 1.8$ no runs yield feasible solutions. Once again, SM-IMPC achieves the largest feasible domain and appears to be essentially unaffected by $\epsilon_{W}$. In particular Fig. 2 suggests that the proposed data-driven approach is competitive with the one based on set-membership identification for moderate levels of process disturbance, while it turns out to be more conservative for higher values of $\epsilon_{W}$.
@@
Fig. 3 shows the trajectories of the closed-loop system (5), (20), and (34) for 100 different realizations of the process disturbance sequence $w(k)\in\mathcal{W}$. It can be observed that the constraints (6) are robustly satisfied and that all trajectories converge to the set $\mathcal{X}_{\infty}$, in accordance with Theorem 2.
@@
In this setting, the average runtime required to solve one instance of problem (34) is 9.5 ms. This indicates that the proposed DD-IMMPC scheme is significantly more computationally efficient than other data-driven predictive control approaches. In particular, the runtime evaluation reported in [13] shows that the methods in [3, 25] require several minutes to solve the corresponding optimization problem.

### 18 Mar 2026 v2

@@ Abstract
This paper presents a new data-driven robust predictive control law, for linear systems affected by unknown-but-bounded process disturbances. A sequence of input-state data is used to construct a suitable uncertainty representation based on interval matrices. Then, the effect of uncertainty along the prediction horizon is bounded through an operator leveraging matrix zonotopes. This yields a tube that is exploited within a variable-horizon optimal control problem, to guarantee robust satisfaction of state and input constraints. The resulting data-driven predictive control scheme is proven to be recursively feasible and practically stable. A numerical example shows that the proposed approach compares favorably to existing methods based on zonotopic tubes.
@@ I Introduction
In this paper, a new data-driven robust predictive control scheme is proposed, for linear discrete-time systems affected by unknown-but-bounded process disturbances. The set of systems compatible with a collection of state-input data is characterized in terms of interval matrices, in two distinct ways: using a data-driven approach and via set-membership identification. Then, the uncertainty propagation along the prediction horizon is approximated by using a suitable operator based on matrix zonotopes. The considered interval matrix uncertainty representation has a twofold benefit: i) it allows one to guarantee recursive feasibility and practical stability of a variable-horizon robust predictive control scheme, without requiring any assumption on the data density; ii) it provides an efficient way to compute a tube bounding the uncertainty propagation, which turns out to be less conservative than that used in [13]. A further contribution is the comparison of the two proposed robust predictive control laws, based respectively on data-driven and set-membership uncertainty representation, showing that the two approaches yield comparable results for process disturbances of moderate size.
The rest of the paper is organized as follows. Section II introduces some preliminary material on matrix operations and set approximations. The problem formulation is presented in Section III. The proposed solution based on interval matrix uncertainty propagation is detailed in Section IV, which also states the closed-loop properties of the control scheme. Section V reports numerical results showing the performance of the proposed approach. Finally, some conclusions are drawn in Section VI.
@@ II Notation and preliminaries
For sets of real matrices $\mathcal{A}$, $\mathcal{B}$, the following operations are considered: $\mathcal{A}\oplus\mathcal{B}=\{A+B|\,\,A\in\mathcal{A},B\in\mathcal{B}\}$, $\mathcal{A}\ominus\mathcal{B}=\{M|\,\,M\oplus\mathcal{B}\subseteq\mathcal{A}\}$, $\mathcal{A}\mathcal{B}=\{AB|\,\,A\in\mathcal{A},B\in\mathcal{B}\}$, $\mathcal{A}^{j}=\{A^{j}|\,\,A\in\mathcal{A}\}$. An interval matrix $\mathcal{I}=\{M|\,\,C-\Delta\leq M\leq C+\Delta\}$, where the operator $\leq$ represents elementwise inequality, is compactly denoted as $\mathcal{I}=C\oplus\left\llbracket\Delta\right\rrbracket$. A matrix zonotope $\mathcal{M}=\Big\{M|M=M_{C}+\sum_{i=1}^{n_{g}}G_{i}\beta_{i},\;|\beta_{i}|\leq 1\Big\}$, where $M_{C}$ is the center and $G_{i}$, $i=1,\dots,n_{g}$, are the generators, is denoted as $\mathcal{M}=\langle M_{C};\,G_{1},\ldots,\,G_{n_{g}}\rangle$.
@@
Given a matrix $M\in\mathbb{R}^{l\times q}$, the set of matrices $\mathcal{E}(M)=\left\{M^{(k)}\right\}_{k=1}^{lq}$, contains all matrices $M^{(k)}\in\mathbb{R}^{l\times q}$ defined as
$M^{(k)}_{ij}=\begin{cases}M_{ij},&\text{if }k=(i-1)q+j,\\ 0,&\text{otherwise},\end{cases}$
@@
If $j$ successive iterations of the operator $\mathbb{T}_{\mathcal{I}}$ are performed on the matrix zonotope $\mathcal{M}$, one obtains the set approximation
@@ III Problem Formulation
Consider the linear time-invariant system
@@
where $x(k)\in\mathbb{R}^{n}$ is the system state at time $k$, $u(k)\in\mathbb{R}^{m}$ is the control input, $w(k)\in\mathbb{R}^{n}$ is the process disturbance and A,BA,\,B are unknown system matrices. It is assumed that $w(k)$ is bounded componentwise, i.e.
@@
Suppose that a sequence of input-state data of length $T$ generated by system (5) is available. These samples are organized into the data matrices
$X_{+}$ $=\left[x(1)\;x(2)\;\cdots\;x(T)\right],$ (6)
$X_{-}$ $=\left[x(0)\;x(1)\;\cdots\;x(T-1)\right],$ (7)
$U_{-}$ $=\left[u(0)\;u(1)\;\cdots\;u(T-1)\right].$ (8)
@@
for any possible realization of $w(k)$, with $\mathcal{X}$, $\mathcal{U}$ being convex polytopic sets.
To proceed, we first describe two different ways to compute interval matrix bounds for the unknown system matrices $A$ and $B$, that are consistent with the collected dataset: one directly based on data matrices (6)-(8), and one exploiting set-membership identification. Then, a general framework is introduced as a basis for the proposed data-driven robust predictive control approach.
@@ III-A Data-driven Interval System Matrices
with $\Delta_{S}\!=\!\overline{w}\,\mathbbm{1}^{\top}|\Phi^{\dagger}|$, where $\mathbbm{1}$ denotes the all-ones vector and $\Phi^{\dagger}$ is the right pseudoinverse of $\Phi$.
@@ III-B Set-membership identification of Interval System Matrices
where $X_{+,i}$ is the $i$-th row of $X_{+}$. Therefore, the feasible set for each $\theta_{i}$ is a bounded polyhedron, and the tightest interval bounds on the entries of $\Theta$ can be computed element-wise by solving $n(n+m)$ linear programs subject to the constraints (21). Once such bounds are derived, an uncertainty model in the form (17) is available and a robust predictive control scheme can be designed on system (18).
@@ III-C Robust Predictive Control
IV Interval-Matrix Robust Predictive Control
@@
$\mathcal{B}^{\mathcal{W}}(j)=\sum_{i=0}^{j-1}\mathcal{I}^{\mathcal{W}}(i)$ (33)
@@
$SS_{k}(j)\subseteq\mathcal{B}_{k}(j),\quad\forall k,j.\vskip 5.69054pt$ (36)
Proof. From (24) one has that $\mathcal{A}_{K}\subseteq\mathcal{I}_{A_{K}}$. Hence, as a consequence of (4) one gets $\mathcal{A}_{K}^{j}\mathcal{I}_{\Delta}\subseteq\mathcal{I}^{\Delta}(j)$ in (32) and $\mathcal{A}_{K}^{j}\mathcal{W}\subseteq\mathcal{I}^{\mathcal{W}}(j)$ in (34). Therefore, the result follows by comparing (30) and (35).∎
@@
Denoting by $N^{*}_{k}$, $\mathbf{v}^{*}(k)$, $\mathbf{z}^{*}(k)$ the optimal solution of problem (37), at each time $k$ the nominal control input in (23) is chosen as
@@
like in standard MPC schemes.
The only difference between problems (37) and (27) is that the constraints (37c)-(37e) are more conservative than the corresponding constraints in problem (27), due to the inclusion (36). This implies that the closed-loop system resulting from the application of the control law (37)-(38) to system (5) will robustly satisfy the constraints (9), whatever are the system matrices $A,B$ and the process disturbance realization $w(k)\in\mathcal{W}$.
It is worth stressing that, while the sets $\mathcal{B}_{k}^{\Delta}(j)$ depend on the nominal state and input sequences $\mathbf{z}_{k}$ and $\mathbf{v}_{k}$, which are optimization variables of problem (37), the interval matrices $\mathcal{I}^{\Delta}(j)$ and the sets $\mathcal{I}^{\mathcal{W}}(j)$, $\mathcal{B}^{\mathcal{W}}(j)$ can be computed offline for all $j$, since they are independent of $\mathbf{z}_{k}$ and $\mathbf{v}_{k}$. This is a key feature of the proposed approach because it allows to significantly limit the computational burden required for the online solution of problem (37).
In order to derive the closed-loop properties of the control scheme (37)-(38), the following assumptions are enforced.
@@
The terminal set $\mathcal{X}_{f}$ in (37e) is a robust positive invariant (RPI) set for the system $x(k+1)=A_{K}x(k)+w(k)$, which means that $A_{K}x+w\in\mathcal{X}_{f}$, $\forall x\in\mathcal{X}_{f}$, $\forall A_{K}\in\mathcal{A}_{K}$ in (24), and $\forall w\in\mathcal{W}$. Moreover, $\mathcal{X}_{f}$ is such that $\mathcal{X}_{f}\subseteq\mathcal{X}$ and $K\mathcal{X}_{f}\subseteq\mathcal{U}$.
@@
Let Assumptions 1-3 be satisfied. Then, the closed-loop trajectories $x(k)$ of system (5), under the control law (23), (37)-(38), reach the set $\mathcal{X}_{f}$ in at most $\lfloor J_{0}^{*}/\gamma\rfloor$+1 steps and converge asymptotically to a set included in $\mathcal{X}_{\infty}=\sum_{i=0}^{\infty}\mathbb{T}_{\mathcal{I}_{A_{K}}}^{i}(\mathcal{W})$.
@@
It is worth remarking that the advantage of the proposed approach stems from the combined use of two distinct features. First, the zonotopic uncertainty propagation along the prediction horizon, through the operator $\mathbb{T}^{j}$, permits to explicitly incorporate the effect of the data-driven uncertainty model (15)-(16) in the robust optimal control problem (37). Secondly, the adoption of the interval-matrix-based tube $\mathcal{B}_{k}(j)$ has a twofold effect: (i) it is key to prove recursive feasibility of the predictive control scheme; (ii) it allows one to cast the constraints of problem (37) in the computationally lightweight form (40).
The adoption of a variable-horizon predictive control scheme is instrumental to achieve recursive feasibility. This makes problem (37) a mixed-integer program with linear constraints and only one integer variable $N_{k}$. Such a problem can be solved exactly by using standard solvers (see, e.g., [16]), or by adopting efficient heuristic procedures [20, 23].
@@ V Numerical results
In this section, numerical results for a double-integrator example are presented to evaluate the performance of two control schemes: i) the Data-Driven Interval Matrix Predictive Control (DD-IMPC) scheme, in which the control law (23), (37)-(38) is applied to the data-driven interval matrices (13); ii) the Set-Membership Interval Matrix Predictive Control (SM-IMPC) scheme, in which the same control law is applied to the uncertainty model (17), estimated via the set-membership procedure in Section III-B. These control strategies are compared to the Tube-Based Zonotopic Predictive Control (TZPC) method introduced in [13]. The simulations are carried out using MATLAB on a laptop equipped with the Apple M4 processor and 16 GB of RAM. The code is written in MATLAB and is publicly available at https://github.com/RenatoQuartullo/Data-Driven-Interval-Matrices-PC.
Consider the discrete-time system (5), in which the true system matrices are
@@
while the input command must satisfy $|u(k)|\leq 2$. The matrix gain $K$ and the terminal set $\mathcal{X}_{f}$ are calculated following the procedures described in [19, Sec. 5.2] and [6, Ch. 10], respectively. In DD-IMPC and SM-IMPC the function $\ell(\cdot)$ in the cost $J_{k}$ is quadratic, i.e., $J_{k}=\gamma N_{k}+\sum_{j=0}^{N_{k}-1}\|v_{k}(j)-Kz_{k}(j)\|_{2}^{2}$ with $\gamma=1$. The variable-horizon problem (37) is solved by enumeration for each horizon length $N_{k}\in\{1,\ldots,10\}$. The true system is used only to generate the data matrices $X_{-},\,X_{+},\,U_{-}$, collecting a trajectory of length $T=50$.
First, we evaluate the feasible domain (FD) of the considered control schemes and its sensitivity to the available data. The FDs are computed as the convex hull of the initial conditions for which the initial optimization problem is feasible. A total of 300 initial conditions are generated by uniformly gridding the admissible state set $\mathcal{X}$. To ensure a fair comparison, the feasible domain of TZPC (which relies on a fixed prediction horizon) is defined as the set of all initial states for which the corresponding optimization problem is feasible for at least one horizon length $N\in\{1,\ldots,10\}$. The FD evaluation is repeated 50 times using independently generated input-state datasets. Table I reports for each method the average area of the FDs, its standard deviation and the number of datasets for which the FD is empty. The results show that DD-IMPC achieves a larger FD on average and exhibits significantly lower sensitivity to the available data compared to TZPC. It is also worth noting that the FD of DD-IMPC is very close to that yielded by SM-IMPC (which is essentially unaffected by the dataset), although the latter exploits an intermediate identification step. Figure 1 illustrates the FDs obtained in one representative run in which TZPC produces a non-empty feasible domain.
@@
To assess the sensitivity of the considered control schemes to the size of the process disturbance, we consider $\mathcal{W}=\epsilon_{W}\left\llbracket\overline{w}\right\rrbracket$, with $\epsilon_{W}\in\{0.2,0.4,\ldots,0.8,2\}$. For each value of $\epsilon_{W}$, the FD is estimated 50 times using independently generated datasets. The average area of the resulting FDs is reported in Fig. 2 (solid lines), along with the minimum and maximum values across the runs (dashed lines), for the different values of $\epsilon_{W}$. It can be observed that, on average, DD-IMPC yields a larger feasible domain than TZPC. Moreover, TZPC exhibits a strong sensitivity to $\epsilon_{W}$, as the area of the FD drops significantly for larger values of the disturbance bound. Moreover, the low data sensitivity of DD-IMPC is confirmed by the small gap between the minimum and maximum area profiles, whereas this effect is much more pronounced for TZPC. In addition, for $\epsilon_{W}\geq 0.8$, TZPC yields an empty feasible region in some runs, while for $\epsilon_{W}\geq 1.8$ all datasets generate an infeasible problem. Once again, SM-IMPC achieves the largest feasible domain and appears to be essentially unaffected by $\epsilon_{W}$. In particular Fig. 2 suggests that DD-IMPC is competitive with SM-IMPC for moderate levels of process disturbance, while it turns out to be more conservative for higher values of $\epsilon_{W}$.
@@
Fig. 3 shows the trajectories of the closed-loop system resulting from one instance of the DD-IMPC scheme, for 100 different realizations of the process disturbance sequence $w(k)\in\mathcal{W}$. It can be observed that the constraints (9) are robustly satisfied and that all trajectories converge to the set $\mathcal{X}_{\infty}$, in accordance with Theorem 2.
@@
In this setting, the average runtime required to solve one instance of problem (37) is 9.5 ms, while that for one iteration of TZPC is 1.6 ms. This is expected, taking into account that TZPC employs a fixed horizon and a rigid tube, while (37) is a variable-horizon problem that explicitly accounts for uncertainty propagation along the prediction horizon. On the other hand, it is remarkable that the runtime of the proposed method is much smaller than that of existing DDPC techniques which incorporate uncertainty propagation within the optimal control problem, such as [3, 25] that require several minutes per iteration (see [13]).